OpenStudy (anonymous):

Factor the following trinomial. x^2 + 25x + 156

OpenStudy (anikate):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac}\ }{ 2a }\]

OpenStudy (anikate):

plug the numbers in and see what u get

OpenStudy (anonymous):

\[-25\pm \frac{ \sqrt{25^2-4(1)(156)} }{ 2(1) }?\] Like this?

OpenStudy (anikate):

yup!! :D AWesome!!

OpenStudy (anikate):

make sure to do 2 equations where in one it is subtracting and the other it is adding

OpenStudy (anonymous):

\[-25\pm \frac{ \sqrt{1} }{2 }\] or \[-25\pm \frac{ 1 }{ 2 }\] Do I get rid of the square root/division sign or no?

OpenStudy (anikate):

just a sec. lemme check

OpenStudy (anikate):

either way is fine

OpenStudy (anikate):

i would prefer the second one just to make it easier and removing as many roots as possible for whole numbers

OpenStudy (anonymous):

now i separate it?

OpenStudy (anikate):

yes do 2 equations, one would be adding and the other will be subtracting

OpenStudy (anikate):

aw darn it! sry for sending you on a goose chase, i just made u find the x values not the factorizin. sry man!!

OpenStudy (anikate):

(x+12)(x+13)

OpenStudy (anikate):